# EarthGPT-X: A Spatial MLLM for Multi-level Multi-Source Remote Sensing Imagery Understanding with Visual Prompting

### v2

@@ V-C1 Evaluation on Optical Data
Referring object classification aims to identify the category labels of objects referred to by visual prompts and textual instructions. To assess the merits of the proposed EarthGPT-X, we report its performance and compare it with the recent vanilla and visual prompting MLLMs on DIOR-RSVG[7]. We adapt DIOR-RSVG by converting referring expressions and their bounding boxes into classification samples, where each target region serves as input and its category label as output. The text instruction is “Please identify the object category of the marked region in the image”. The vanilla MLLMs do not support drawing visual prompts on the images, therefore, the coordinates contained in the text instruction serve as the referring regions. The vanilla MLLMs used for comparison include natural-domain models such as Sphinx [89], InternVL2.5-1B[38], InternVL3-1B[91], as well as RS-domain models including GeoChat [87] and EarthGPT [8]. For visual prompting MLLMs, we select the most recent representatives, namely the natural-domain models Sphinx-V [77] and Vip-LLava [18], together with the RS-domain model EarthMarker [15]. We conduct experiments under three levels of visual prompting (point, box, and free-form) to provide a fair and comprehensive comparison across models. Following [77, 1], for point prompts, we randomly select a point within the ground-truth object that is also close to its boundary. For box prompts, we use the ground-truth bounding box provided by DIOR-RSVG. For free-form prompts, we randomly perturb the center positions and zoom the box prompts to simulate. As shown in Tab. IV, EarthGPT-X achieves optimal results in S-IoU under box prompts. We also observe that the box-level results are better than the point and free-form prompt ones.
@@ V-D Image Captioning
To evaluate the image captioning capabilities, we adopt the NWPU-Captions [94] dataset to benchmark EarthGPT-X against other expert models under supervised settings. During evaluation, we use the instruction “Please provide a brief caption of each marked region in the image” alongside a full-image box $0,0,\mathrm{width},\mathrm{height}]$ as the visual prompt to guide the captioning process. As reported in Tab. VI, EarthGPT-X consistently outperforms existing expert methods across all evaluation metrics. These results indicate that by treating the entire image as a unified region and leveraging multi-domain training on both natural and RS data, EarthGPT-X is capable of generating accurate and semantically rich descriptions for RS scenes.
@@ V-E1 Evaluation on Optical Data
For the brief region captioning on optical data, the test set of OPT-RSVG [98] is employed for zero-shot evaluation. We elaborately choose both natural and RS-specific MLLMs for performance comparison. Natural scene MLLMs contains Qwen-VL [9], InternVL2.5-1B[38], InternVL3-1B[91], Sphinx [89], Sphinx-V [77], Vip-LLava [18] and GLaMM [18]. Popular RS-specific MLLMs GeoChat [87] and EarthGPT [8] are selected. The text instruction for the region captioning task is designed as “Please provide a brief caption of each marked region in the image”. As shown in Tab. VII, the proposed EarthGPT-X consistently outperforms previous works across all evaluation metrics. Notably, EarthGPT-X achieves significant gains over Sphinx-V, despite the latter being trained on large-scale RS datasets. This highlights the superior performance and stronger generalization ability of EarthGPT-X in multi-source RS data understanding. Notably, using boxes as visual prompts can yield the best performance on region captioning ability.
@@ V-E2 Evaluation on SAR and Infrared Data
As displayed in Tab. VIII, for the multi-source brief captioning task, we evaluate EarthGPT-X on two representative datasets: SAR-Ship [99] for supervised evaluation and HIT UAV [100] under a zero-shot setting. These datasets enable us to validate the generalization ability of EarthGPT-X across different modalities. For comparison, we include two representative general-purpose MLLMs, GPT-4o and InternVL2 [90]. We further reference the first RS MLLM, GeoChat [9], and the latest multi-source understanding model, EarthDial [23], as strong domain-related baselines. Experimental results demonstrate that EarthGPT-X achieves superior overall performance on multi-source datasets.
@@ V-F Ablation Study
Multi-scale + Sub-images MVE 98.60 97.83
@@ V-F1 Different Visual Encoders
To show the effectiveness of the proposed hybrid MVE module, we ablate on the performance of our EarthGPT-X when using different visual encoders. Specifically, this ablation uses CLIP-ConvNeXt-L, CLIP RN50×16, and DINOv2 ViT-L/14 for comparison. The results are shown in Tab. IX, it is clear that the setting using the ConvNeXt-L and DINOv2 ViT-L/14 as the mixed visual experts achieves the most superior performance on scene classification, region captioning, and referring object classification tasks. These results demonstrate that utilizing a network architecture with complementary functions has deepened the understanding of local information and contextual dependencies on RS imagery.
@@ V-F2 Sub-images Decomposition and Multi-scale
In the visual feature extraction stage, we strengthen feature integration by leveraging multi-scale features and high-resolution sub-images for richer RS image comprehension. Ablation results in Tab. X show that this design significantly outperforms single-scale settings on the DIOR-RSVG dataset, confirming that capturing finer details through multiple scales and sub-images is crucial for accurate interpretation.
@@ VI-D Multiple Visual Prompts
To evaluate the adaptability of EarthGPT-X under different prompting shapes, we investigate point prompt, box prompt, and free-form prompt. These prompts vary in the way visual regions are highlighted to the model, thereby influencing the style and granularity of responses. i) Similarities: Across all prompt types, EarthGPT-X consistently grounds the responses to the marked regions, ensuring semantic coherence and correct alignment with the visual input. ii) Differences: As shown in the Fig. 8, under the point prompt, EarthGPT-X generates a caption emphasizing the spatial relation, which reflects its local context-awareness. Under the box prompt, the model highlights the absolute spatial location, showing a region-level grounding behavior. In contrast, the free-form prompt yields a more generic description, focusing on object-level attributes (Note that these examples are qualitative illustrations, not unique or deterministic results, since different prompt placements may lead to varied outputs).
@@ VII Conclusion
In this paper, we propose EarthGPT-X, a flexible and versatile MLLM specifically designed for understanding RS imagery from multiple sources. Technically, the hybrid signal mutual understanding mechanism is developed to enhance the interplay between dense visual features, sparse free-form visual prompts, and text instructions, facilitating a multi-grained understanding of complex RS imagery. Furthermore, a novel cross-domain one-stage fusion training strategy is proposed to generalize general-domain knowledge into RS multi-source scenarios. Through the designed visual prompt learning framework and training strategy, EarthGPT-X achieves multi-level visual understanding at the image, region, and pixel levels, etc., allowing for precise and intelligent analyses in real-world applications. In the future, we aim to expand EarthGPT-X towards a wider range of tasks and more advanced visual reasoning capabilities across diverse modalities, ultimately strengthening its versatility and effectiveness in comprehensive RS imagery analysis.

### v3

@@ V-C1 Evaluation on Optical Data
Referring object classification aims to identify the category labels of objects referred to by visual prompts and textual instructions. To assess the merits of the proposed EarthGPT-X, we report its performance and compare it with the recent vanilla and visual prompting MLLMs on DIOR-RSVG[2]. We adapt DIOR-RSVG by converting referring expressions and their bounding boxes into classification samples, where each target region serves as input and its category label as output. The text instruction is “Please identify the object category of the marked region in the image”. The vanilla MLLMs do not support drawing visual prompts on the images, therefore, the coordinates contained in the text instruction serve as the referring regions. The vanilla MLLMs used for comparison include natural-domain models such as Sphinx [90], InternVL2.5[37], InternVL3[92], as well as RS-domain models including GeoChat [88] and EarthGPT [7]. For visual prompting MLLMs, we select the most recent representatives, namely the natural-domain models Sphinx-V [76] and Vip-LLava [17], together with the RS-domain model EarthMarker [14]. We conduct experiments under three levels of visual prompting (point, box, and free-form) to provide a fair and comprehensive comparison across models. Following [76, 1], for point prompts, we randomly select a point within the ground-truth object that is also close to its boundary. For box prompts, we use the ground-truth bounding box provided by DIOR-RSVG. For free-form prompts, we randomly perturb the center positions and zoom the box prompts to simulate. As shown in Tab. IV, EarthGPT-X achieves optimal results in S-IoU under box prompts. We also observe that the box-level results are better than the point and free-form prompt ones.
@@ V-D Image Captioning
To evaluate the image captioning capabilities, we adopt the NWPU-Captions [95] dataset to benchmark EarthGPT-X against other expert models under supervised settings. During evaluation, we use the instruction “Please provide a brief caption of each marked region in the image” alongside a full-image box $0,0,\mathrm{width},\mathrm{height}]$ as the visual prompt to guide the captioning process. As shown in Table VI, EarthGPT-X outperforms most expert-designed models but remains inferior to task-specific architectures such as Aware-Transformer [97]. We note that the performance gap in image captioning is a common limitation of MLLMs, which aim to balance multiple capabilities rather than optimizing for a single task.
@@ V-E1 Evaluation on Optical Data
For the brief region captioning on optical data, the test set of OPT-RSVG [101] is employed for zero-shot evaluation. We elaborately choose both natural and RS-specific MLLMs for performance comparison. Natural scene MLLMs contain Qwen-VL [8], InternVL2.5[37], InternVL3[92], Sphinx [90], Sphinx-V [76], Vip-LLava [17] and GLaMM [17]. Popular RS-specific MLLMs GeoChat [88] and EarthGPT [7] are selected. The text instruction for the region captioning task is designed as “Please provide a brief caption of each marked region in the image”. As shown in Tab. VII, the proposed EarthGPT-X consistently outperforms previous works across all evaluation metrics. Notably, EarthGPT-X achieves significant gains over Sphinx-V, despite the latter being trained on large-scale RS datasets. This highlights the superior performance and stronger generalization ability of EarthGPT-X in multi-source RS data understanding. Notably, using boxes as visual prompts can yield the best performance on region captioning ability.
@@ V-E2 Evaluation on SAR and Infrared Data
As displayed in Tab. VIII, for the multi-source brief captioning task, we evaluate EarthGPT-X on two representative datasets: SAR-Ship [67] for supervised evaluation and HIT UAV [102] under a zero-shot setting. These datasets enable us to validate the generalization ability of EarthGPT-X across different modalities. For comparison, we include two representative general-purpose MLLMs, GPT-4o from OpenAI and InternVL2 [91] from OpenGVLab. We further reference the first RS MLLM, GeoChat [8], and the latest multi-source understanding model, EarthDial [22], as strong domain-related baselines. Experimental results demonstrate that EarthGPT-X achieves superior overall performance on multi-source datasets.
@@ V-F Ablation Study
Multi-scale + Sub-images MVE 98.79 98.03
Notes: Results are reported under box-prompt settings.
@@ V-F1 Different Visual Encoders
To show the effectiveness of the proposed hybrid MVE module, we ablate on the performance of our EarthGPT-X when using different visual encoders. Specifically, this ablation uses CLIP-ConvNeXt-L, CLIP RN50×16, and DINOv2 ViT-L/14 for comparison. The results are shown in Tab. IX, it is clear that the setting using the CLIP-ConvNeXt-L and DINOv2 ViT-L/14 as the mixed visual experts achieves the most superior performance on scene classification, region captioning, and referring object classification tasks. These results demonstrate that utilizing a network architecture with complementary functions has deepened the understanding of local information and contextual dependencies on RS imagery.
@@ V-F2 Sub-images Decomposition and Multi-scale
In the visual feature extraction stage, we strengthen feature integration by leveraging multi-scale features and high-resolution sub-images for richer RS image comprehension. Ablation results in Tab. X show that this design significantly outperforms single-scale settings under the referring object classification task, confirming that capturing finer details through multiple scales and sub-images is crucial for accurate interpretation.
@@ VI-D Multiple Visual Prompts
To evaluate the adaptability of EarthGPT-X under different prompting shapes, we investigate point prompt, box prompt, and free-form prompt. These prompts vary in the way visual regions are highlighted to the model, thereby influencing the style and granularity of responses. i) Similarities: Across all prompt types, EarthGPT-X consistently grounds the responses to the marked regions, ensuring semantic coherence and correct alignment with the visual input. ii) Differences: As shown in Fig. 8, under the point prompt, EarthGPT-X generates a caption emphasizing the spatial relation, which reflects its local context-awareness. Under the box prompt, the model highlights the absolute spatial location, showing a region-level grounding behavior. In contrast, the free-form prompt yields a more generic description, focusing on object-level attributes (Note that these examples are qualitative illustrations, not unique or deterministic results, since different prompt placements may lead to varied outputs).
@@ VII Conclusion
In this paper, we propose EarthGPT-X, a flexible and versatile MLLM specifically designed for understanding RS imagery from multiple sources. Technically, the hybrid signal mutual understanding mechanism is developed to enhance the interplay between dense visual features, sparse free-form visual prompts, and text instructions, facilitating a multi-grained understanding of complex RS imagery. Furthermore, a novel cross-domain one-stage fusion training strategy is proposed to generalize general-domain knowledge into RS multi-source scenarios. Through the designed visual prompt learning framework and training strategy, EarthGPT-X achieves multi-level visual understanding at the image, region, and pixel levels, etc., allowing for precise and intelligent analyses in real-world applications. In the future, we plan to enhance the image captioning capability of EarthGPT-X through task-aware fine-tuning and adaptive prompt optimization. Furthermore, we aim to expand EarthGPT-X towards a wider range of tasks and more advanced visual reasoning capabilities across diverse modalities, ultimately strengthening its versatility and effectiveness in comprehensive RS imagery analysis.